OpenStudy (anonymous):

Given that theta = sec^-1(2.6), find the exact values of sin theta, cos theta, tan theta, cot theta, and csc theta.

OpenStudy (anonymous):

\[\Theta=\sec ^{-1}2.6\]

myininaya (myininaya):

2.6=26/10=13/5 \[\sec(\theta)=\frac{hyp}{adj}=\frac{13}{5}\]  we can find the other side by using Pythagorean thm